Can you please help me solve this problem?
The area of a circle is 268 square inches. What is the radius to the nearest inch?
Thank you!

Answer by rfer(16322)

A=3.14r^2
268=3.14r^2
268/3.14=r^2
85.35=r^2
r=sqrt 85.35
r=9 in